OpenStudy (anonymous):

Use multiplying by 1 to find an expression equivalent to the fraction of 20 over 17 with a denominator of 34x

OpenStudy (anonymous):

20/17=a/34x so multiply top and bottom of left hand side by 2 (that is your 2/2 = 1) to give 40/34 =a/34x. Now a=40x.